OpenStudy (anonymous):

what is the simplified form of 1/10xy * 5x^2/7y ?

OpenStudy (campbell_st):

do the multiplication 1st then simplify \[\frac{1}{10xy} \times \frac{5x^2}{7y} = \frac{5x^2}{70xy^2}\] now eliminate common factors \[\frac{5}{70} = \frac{1}{14} \] and \[\frac{x^2}{x} = \frac{x}{1}\] so the simplification is \[\frac{x}{14y^2}\]

OpenStudy (anonymous):

thanksss!!!